Circular No. 5690

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 C. Bartolini, A. Guarnieri, and A. Piccioni, Bologna University, com-

municate: “During monitoring of GRO J0422+32 with the Bologna Observatory 1.5-m telescope (+ fast double-head photometer + V filter; integration time 1 s) on 1992 Dec. 30.0210 UT, the star dropped abruptly by > 3 magnitudes in 20 s. For an hour and a half, the star was not detectable, and the monitoring then was stopped. We have no plausible explanations that attribute this fact to observational errors. CCD observations in B, V, R, and I on 1993 Jan. 4.88 and 5.80 show that the star is again visible.”

Visual magnitude estimates (cf. IAUC 5613): 1992 Sept. 23.16 UT, 13.0 (P. Schmeer, Bischmisheim, Germany); 27.15, 13.1 (Schmeer); Oct. 19.12, 13.2 (Schmeer); 20.08, 13.2 (Schmeer); Dec. 16.77, 14.5: (J.-C. Merlin, Le Creusot, France, 0.40-m reflector); 29.92, 15.0: (Merlin).

PKS 1406–076 The Compton Observatory EGRET Team (B. L. Dingus, D. L. Bertsch,

C. E. Fichtel, R. C. Hartman, S. D. Hunter, P. W. Kwok, J. R. Mattox, P. Sreekumar, and D. J. Thompson, Goddard Space Flight Center, NASA; D. A. Kniffen, Hampden-Sydney College; J. Chiang, J. Fierro, Y. C. Lin, P. F. Michelson, and P. L. Nolan, Stanford University; E. Schneid, Grumman Aerospace Corporation; and K. Brazier, G. Kanbach, H. A. Mayer-Hasselwander, and C. von Montigny, Max-Planck-Institut für Extraterrestrische Physik) communicates: “We report the tentative identification of the quasar PKS 1406–076 in the γ-ray energy range above 100 MeV for the observation interval 1992 Dec. 22 through 1993 Jan. 3. This quasar is 0^{o}.2 from the γ-ray position and is the only active galactic nucleus within the 0^{o}.5 68-percent confidence contour. EGRET will continue to view this region of the sky until Feb. 2, and a planned observation from July 8 to July 13 will view this source.”

COMET OHSHITA (1992a_{1}) Total visual magnitude estimates: 1992 Dec. 23.52 UT, 12.2 (A. Hale,

Las Cruces, NM, 0.41-m reflector); 24.52, 11.3 (C. S. Morris, Whitaker Peak, CA, 0.26-m reflector); 27.07, 11.9 (H. Mikuz, Ljubljana, Slovenia, 0.20-m reflector); 30.05, 12.6 (Mikuz); 1993 Jan. 3.55, 11.9 (Morris).

1993 January 7 Daniel W. E. Green